OpenStudy (anonymous):

Convert the polar equation to rectangular equation r=1/(1+sin theta) my answer is y^2(x+1)=r or (y^2(x+1))^2=x^2+y^2 I need somebody to check it for me :D

OpenStudy (anonymous):

@Kainui yes you are here help plz

OpenStudy (perl):

did you cross multipl

OpenStudy (anonymous):

nop

OpenStudy (perl):

actually lets use the conversion factors x = r cos t y = r sin t x^2 + y^2 = r^2 so we have y/r = sin t

OpenStudy (anonymous):

want me to draw how i got the answer?

OpenStudy (perl):

r = 1 / ( 1 + y/r ) now cross multiply

OpenStudy (perl):

r ( 1 + y/r) = 1 r*1 + r* (y/r) = 1 r + y = 1 r = 1 - y now square both sides

OpenStudy (perl):

r^2 = ( 1 - y ) ^2 x^2 + y^2 = ( 1 - y)^2

OpenStudy (perl):

what did you get?

OpenStudy (anonymous):

let me draw it

OpenStudy (anonymous):